OpenStudy (anonymous):

An aid package is to be dropped from a plane flying at an elevation of 300m to stranded tourist. if the plane is flying at 65ms-1 calculate the distance before tourist the package must be released in order to reach tourist.

OpenStudy (anonymous):

first find how long it takes the package to hit the ground 300 = (1/2) (9.8) t^2 t = 7.82s d = vt = 65 * 7.82 = 508.3m <--